OpenStudy (anonymous):

Factor completely: 3x2 - x - 4

OpenStudy (anonymous):

(x+1)(3x−4)

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

I just used this: https://www.mathway.com/ Much easier and faster than solving it on paper.

OpenStudy (anonymous):

(3x - 1)(x + 4) (3x + 4)(x - 1) (3x - 2)(x + 2) (3x - 4)(x + 1) These are the answers I can choose from though

OpenStudy (anonymous):

So it's just the other way around?

OpenStudy (anonymous):

The answer's D :)

OpenStudy (anonymous):

Yeah the other way around. C:

OpenStudy (anonymous):

thx @maria195 wow such a cheater

OpenStudy (anonymous):

@amistre64 can you explain it to me

OpenStudy (anonymous):

@camerondoherty

OpenStudy (anonymous):

@yamyam70

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

could you guys explain why is this the answer

OpenStudy (anonymous):

i dont understand how can this be the answer

OpenStudy (amistre64):

what does it mean to be a factor?

OpenStudy (yamyam70):

@Law&Order try using the FOIL method, to check

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

resolve or be resolvable into factors

OpenStudy (amistre64):

if i ask you what are the factors of 25, what am i asking for?

OpenStudy (anonymous):

the factors are 5*5

OpenStudy (amistre64):

so, what you have done is to split it up into multiplication, right?

OpenStudy (anonymous):

yes, ok im getting it

OpenStudy (camerondoherty):

Step 1 : Simplify 3x2-x - 4 Trying to factor by splitting the middle term 1.1 Factoring 3x2-x-4 The first term is, 3x2 its coefficient is 3 . The middle term is, -x its coefficient is -1 . The last term, "the constant", is -4 Step-1 : Multiply the coefficient of the first term by the constant 3 • -4 = -12 Step-2 : Find two factors of -12 whose sum equals the coefficient of the middle term, which is -1 . -12 + 1 = -11 -6 + 2 = -4 -4 + 3 = -1 That's it Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, -4 and 3 3x2 - 4x + 3x - 4 Step-4 : Add up the first 2 terms, pulling out like factors : x • (3x-4) Add up the last 2 terms, pulling out common factors : 1 • (3x-4) Now add up the four terms of step 3 : (x+1) • (3x-4) Which is the desired factorization Final result : (3x - 4) • (x + 1)

OpenStudy (anonymous):

\(\huge\ddot\smile\)

OpenStudy (anonymous):

is that tiger algebra

OpenStudy (amistre64):

when we factor something like: 3x2 - x - 4 we want to turn it into a multiplication setup (this) (that)

OpenStudy (camerondoherty):

yup :)

OpenStudy (anonymous):

im sorry but i dont understand

OpenStudy (camerondoherty):

lol

OpenStudy (anonymous):

oh @amistre64

OpenStudy (anonymous):

so we split it up

OpenStudy (anonymous):

but how @amistre64

OpenStudy (amistre64):

there are a few methods to approach it, but all of them are a bit involved and assume you have a basic knowledge of algebra to play with

OpenStudy (anonymous):

yes i think i know

OpenStudy (anonymous):

but there is a way that i did it but it gave me the wrong answer 3x^2-x-4 then i split it to x(3x+3)-2(2x+2)

OpenStudy (amistre64):

lets assume it factors into something like: (ax+b) (cx+d); we can then expand this out to compare parts with ax+b cx+d ------- acx^2 + bcx +adx +bd -------------------- acx^2 +(ad+bc)x + bd which may or maynot be practical, but it helps see the mechanics going on

OpenStudy (anonymous):

\(\color{green}{\huge\cal 3x^2-x-4}\) \(\color{green}{\huge\cal x(3x+3)-2(2x+2)}\) but you cant make it simplifyed because the two parenthesis isnt the same

OpenStudy (anonymous):

mm hm

OpenStudy (anonymous):

*nods head*

OpenStudy (amistre64):

by comparison of like terms we have: 3 x^2 -1 x - 4 ac x^2 +(ad+bc)x + bd ac = 3; let c = 3/a bd = -4; let d=-4/b ad + bc = -1 -4a/b + 3b/a = -1 -4a + 3b^2/a = -b -4a^2 + 3b^2 = -ab so yeah, this approach doesnt seem practical to me.

OpenStudy (amistre64):

my own prefered method, assuming at factors into rationals is: multiply first and last 3 x^2 - x - 4 ; 3(-4) = -12 since its negative we have opposite signs (x+ p/3) (x- q/3) will be the setup, p and q are factors of 12 since the middle term is a negative value, the bigger term is the negative 4,3 are factors of 12, and -4+3 = -1 sooo (x+ 3/3) (x- 4/3) , simplify (x+1) (x- 4/3) , if theres still a fraction, put the bottom back in front (x+1) (3x- 4)

OpenStudy (amistre64):

thats my take on it, since i am no good at factor by grouping :)